OpenStudy (anonymous):

Divide and simplify. z^2+8z/z^2+4z-32\[\div\]z/z+4

OpenStudy (anonymous):

\[\frac{z+4}{z-4}\]

OpenStudy (aroub):

z^2 +8z z --------- ÷ -------- z^2+4z-32 z+4 = z(z+8) z+4 -------- * ------- (z-4)(z+8) z = z z+4 ---- * ------ z-4 z = z+4 ----- z-4 \[z \neq 4\] & \[z \neq -8 \] & \[z \neq 0 \]